OpenStudy (anonymous):

Can someone help with a complex number problem in euler's form and using de moivre's theorem? Medals given

OpenStudy (astrophysics):

Hey, are you there? What's your question @chau88

OpenStudy (astrophysics):

Ooh interesting, I know that \[e^{i \theta} = \cos \theta + i \sin \theta\] so you will have to apply this

OpenStudy (anonymous):

This question, for (a), (b) and (c) I have queries. (a) I got, w^2=e^4ipi/5, w^3=e^6ipi/5; w^4=e^8ipi/5

OpenStudy (astrophysics):

That looks decent, let me see

OpenStudy (anonymous):

but my answer is different from the given answer, which is w^4=e^4ipi/5, w^3=e^-4pi/5i; w^4=e^-2pi/5 i

OpenStudy (astrophysics):

\[\huge w^2 = \left( e^{2i \pi/5} \right)^2 = e^{4i \pi/5}\] right?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

\[\huge w^3 = \left( e^{2i \pi/5} \right)^3 = e^{-4i \pi/5}\] this is for cubed is it

OpenStudy (anonymous):

wait. why?

OpenStudy (astrophysics):

That's what you have given for the answer in the book

OpenStudy (anonymous):

did you use the anticlockwise?

OpenStudy (anonymous):

because if you multiply it, you should get 6ipi/5

OpenStudy (anonymous):

Ah, I see. Thanks. @Astrophysics , what about (b), wo far I have e^2ipi/5 + e^4i pi/5+ e^-4ipi/5+e^-2ipi/5. can I group them into 2cos2theta+2 cos 4theta?

OpenStudy (anonymous):

How do I get -1?

OpenStudy (astrophysics):

@ganeshie8

OpenStudy (astrophysics):

I think you can use Euler's formula for b, but I'm not sure

OpenStudy (astrophysics):

identity*

OpenStudy (anonymous):

is it z^n+1/z^n=2cos ntheta? I used that, but I got 12 cos^2 theta -8 and not -1 :(

OpenStudy (anonymous):

Note that\[1, \omega, \omega^2, \omega^3, \omega^4 \]are the roots of the equation\[x^5 - 1= 0\]The sum of roots of this equation is \(0\).

ganeshie8 (ganeshie8):

Nice!

OpenStudy (anonymous):

@Beauregard , how can you tell that this is the fifth root of unity?

ganeshie8 (ganeshie8):